1.Mauro Products distributes a single product, a woven basket whose selling price is $20 per unit and whose variable expense is2.Lin Corporation has a single product whose selling price is $140 per unit and whose variable expense is $70 per unit. The com
Answer #1

1.

1. Breakeven point in unit sales 2,900 baskets
2. Breakeven point in dollar sales $58,000
3. Breakeven point in unit sales 3,020 baskets
Breakeven point in dollar sales $60,400

Unit contribution margin = $20-$15 (selling price per unit-variable cost per unit)

=$5

a. Breakeven point in unit sales= Monthly fixed expense/Unit contribution margin

=$14,500/5

=2900 units

b.Breakeven point in dollar sales= Breakeven units* selling price

= 2900*20

=$58,000

c. Breakeven point in unit sales if the fixed expense increases by $600= Monthly fixed expense/Unit contribution margin

   =$14,500+$600/5

   =15,100/5

   =3,020 baskets

d. Breakeven point in dollar sales if the fixed expense increases by $600= Breakeven units* selling price

=3020*20

=$60,400

2.

1. Unit sales to attain target profit 565 units
2. Dollar sales to attain target profit $81,340

Contribution margin per unit= $140-$70 = $70

1. Unit sales to attain target profit= (Fixed cost+ target profit)/contribution margin per unit

= ($32,500+$7,050)/ $70

=$39,550/$70

=565 units

2.Dollar sales to attain target profit= [(Fixed cost+ target profit)/contribution margin per unit]* selling price per unit

=[($32,500+$8,200)/$70]* $140

($40,700/$70)* $140

581 units*$140

=$81,340